# Hybrid-Task Meta-Learning: A Graph Neural Network Approach for Scalable and Transferable Bandwidth Allocation

### v2

@@ V-C2 Meta-Testing with Different System Parameters
In Fig. 8, we evaluate the initial performance and sample efficiency with different $r_{\tau}^{S,\mathcal{I}}$ in support sets and query sets in meta-training. Specifically, we set $r_{\tau}^{S,\mathcal{I}}$ to $10$ Mbps and $1$ Mbps in meta-training in Figs. 8(a) and 8(b), respectively. In Fig. 8(c), $r_{\tau}^{S,\mathcal{I}}$ is randomly selected from the set $\{1,\cdots,10\}$ Mbps in meta-training. In meta-testing, we increase $r_{\tau}^{S,\mathcal{I}}$ from $1$ Mbps to $10$ Mbps. The results in Figs. 8(a) and 8(b) indicate that the gaps between zero-shot learning (with 00 training epochs in meta-testing) and the optimal solution increase with the difference between $r_{\tau}^{S,\mathcal{I}}$ in meta-training and $r_{\tau}^{S,\mathcal{I}}$ in meta-testing. To increase the generalization ability, we can increase the diversity of tasks in meta-training as shown in Fig. 8(c). In this way, our GNN is near-optimal with zero-shot learning.
@@
In Fig. 9, we validated the generalization ability of our GNN with dynamic bandwidth $W_{\tau}^{S,\mathcal{I}}$. In meta-training, $W_{\tau}^{S,\mathcal{I}}$ is randomly selecting from the set $\{10,\cdots,100\}$ MHz. In meta-testing, we increase $W_{\tau}^{S,\mathcal{I}}$ from $10$ to $100$ MHz. The results in Fig. 9 show that our GNN is near-optimal with different values of $W_{\tau}^{S,\mathcal{I}}$. In Fig. 10, we further validate the generalization ability of our GNN with different numbers of users. In meta-training, the number of total users is randomly selected, $U_{\tau}^{S,\mathcal{I}}\in\{10,11,...,30\}$. In meta-testing, we increase the number of total users from $5$ to $50$. The results in Fig. 10 show that the proposed HML can obtain a GNN that has strong generalization ability with different numbers of users. The gap between the GNN and the optimal policy increases slightly with $U_{\tau}^{S,\mathcal{I}}$. This is because the scale of the problem increases with $U_{\tau}^{S,\mathcal{I}}$, and it is more difficult to learn the bandwidth allocation policy of a large-scale problem compared with that of a small-scale problem.
@@ VI Conclusion
In this paper, we developed an HML approach to train a GNN-based scalable bandwidth allocation policy that can generalize well in various communication scenarios, including different number of users, wireless channels, QoS requirements, and bandwidth. The main idea is to train the initial parameters of the GNN with various tasks in meta-training, and then fine-tune the parameters with a few samples in meta-testing. Simulation results showed that the performance gap between the GNN and the optimal policy obtained by an iterative algorithm is less than $5$% in most of the cases. For unseen communication scenarios, the GNN can converge in $10$ to $30$ training epochs, which are much faster than the existing benchmarks. Our approach can be extended beyond bandwidth allocation, such as power allocation, precoding, and repetitions. Nevertheless, the featuring engineering and the structure of GNN in other scenarios deserve further investigation.
@@ Appendix A Proof of Concavity for Secrecy Rate in Long Blocklength Regimes
$\begin{split}\frac{\partial r_{k}^{S,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})}{\partial{w_{\tau,k}^{D,\mathcal{I}}}}=&\frac{\partial\left(r_{k}^{D,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})-r_{k}^{e,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})\right)}{\partial w_{k}}\\ =&\frac{-\zeta_{k}+\ln\left(1+\frac{\zeta_{k}}{w_{\tau,k}^{D,\mathcal{I}}}\right)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})}\\ &-\frac{-\zeta_{k}^{e}+\ln\left(1+\frac{\zeta_{k}^{e}}{w_{\tau,k}^{D,\mathcal{I}}}\right)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})}\\ =&\frac{\ln\left(\frac{w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}}{w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e}}\right)}{\ln(2)}\\ &+\frac{(\zeta_{k}^{e}-\zeta_{k})w_{\tau,k}^{D,\mathcal{I}}}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})},\end{split}$ (18)
@@
$\begin{split}\frac{\partial^{2}r_{k}^{S,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})}{\partial{w_{\tau,k}^{D,\mathcal{I}}}^{2}}=&\frac{\partial}{\partial w_{\tau,k}^{D,\mathcal{I}}}\left(\frac{\partial r_{k}^{S,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})}{\partial{w_{\tau,k}^{D,\mathcal{I}}}}\right)\\ =&\dfrac{(\zeta_{k}^{e}-\zeta_{k})\left((\zeta_{k}^{e}+\zeta_{k})w_{\tau,k}^{D,\mathcal{I}}+2\zeta_{k}^{e}\zeta_{k}\right)}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})^{2}(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})^{2}}.\end{split}$ (19)

### v3

@@ V-C2 Meta-Testing with Different System Parameters
In Fig. 8, we evaluate the initial performance and sample efficiency with different $r_{\tau}^{S,\mathcal{I}}$ in support sets and query sets in meta-training. Specifically, we set $r_{\tau}^{S,\mathcal{I}}$ to $10$ Mbps and $1$ Mbps in meta-training in Figs. 8(a) and 8(b), respectively. In Fig. 8(c), $r_{\tau}^{S,\mathcal{I}}$ is randomly selected from the set $\{1,\cdots,10\}$ Mbps in meta-training. In meta-testing, we increase $r_{\tau}^{S,\mathcal{I}}$ from $1$ Mbps to $10$ Mbps. The results in Figs. 8(a) and 8(b) indicate that the gaps between zero-shot learning (with 0 training epochs in meta-testing) and the optimal solution increase with the difference between $r_{\tau}^{S,\mathcal{I}}$ in meta-training and $r_{\tau}^{S,\mathcal{I}}$ in meta-testing. To increase the generalization ability, we can increase the diversity of tasks in meta-training as shown in Fig. 8(c). In this way, our GNN is near-optimal with zero-shot learning.
@@
In Fig. 9, we validated the generalization ability of our HML with dynamic bandwidth $W_{\tau}^{S,\mathcal{I}}$. In meta-training, $W_{\tau}^{S,\mathcal{I}}$ is randomly selecting from the set $\{10,\cdots,100\}$ MHz. In meta-testing, we increase $W_{\tau}^{S,\mathcal{I}}$ from $10$ to $100$ MHz. The results in Fig. 9 show that our GNN is near-optimal with different values of $W_{\tau}^{S,\mathcal{I}}$.
In Fig. 10, we further validate the generalization ability of our GNN with different numbers of users. In meta-training, the number of total users is randomly selected, $U_{\tau}^{S,\mathcal{I}}\in\{10,11,...,30\}$. In meta-testing, we increase the number of total users from $5$ to $50$. The results in Fig. 10 show that the proposed HML can obtain a GNN that has strong generalization ability with different numbers of users, which is consistent with our analysis in Section IV-C, indicating that parameters of the user vertices are shared and reused, and the trained GNN processes the aggregated information of the scheduled users. This figure also validates that the trained GNN can effectively be applied in networks with user numbers either smaller or larger than the number of users in the meta-training set. The gap between the GNN and the optimal policy increases slightly with $U_{\tau}^{S,\mathcal{I}}$. This is because the scale of the problem increases with $U_{\tau}^{S,\mathcal{I}}$, and it is more difficult to learn the bandwidth allocation policy of a large-scale problem compared with that of a small-scale problem.
Figure 11: Meta-testing with underestimated eavesdropper channel gain, where $r_{\tau}^{S,\mathcal{I}}=10$ Mbps, $W_{\tau}^{S,\mathcal{I}}=100$ MHz, and $U_{\tau}^{S,\mathcal{I}}=10$.
In Fig. 11, we present simulation results assessing the robustness of our HML algorithm when eavesdropper channel gain is underestimated. This figure reveals a decrease in secrecy rate as the degree of underestimation in the eavesdropper channel gain increases, aligning with the behavior predicted by eq. (8). We can also observe from this figure that the performance gap between the optimal iterative algorithm and our HML increases with the value of underestimation of the eavesdropper channel gain. Thus, there is a trade-off between the eavesdropper channel’s accuracy and the choosing of the bandwidth allocation algorithms. Specifically, the optimality gap of our HML algorithm is small when the eavesdropper channel uncertainty is small, so our HML algorithm could save computation time with a tolerable performance loss. However, this optimality gap increases with the eavesdropper channel uncertainty, so we need to carefully evaluate the accuracy and the computation time when choosing the bandwidth allocation algorithm. This trade-off deserves further examination in the design and deployment of secure communication systems, particularly in real-world scenarios where perfect knowledge of the eavesdropper channel may be unattainable and where bandwidth allocation requires high computational complexity.
TABLE IV: Average Computation Time (millisecond)
Scenario HML meta-training (one epoch) HML inference (one sample) Iterative   (one sample)
Fig. 5(a) 72.61 3.11 641.62
Fig. 5(b) 73.03 3.12 991.47
Fig. 6(a) 103.81 3.25 659.54
Fig. 6(b) 92.16 3.52 2325.60
Fig. 7(a) 78.28 3.62 5392.97
Fig. 7(b) 74.10 3.11 8266.69
V-D Numerical Results of Computation Time
Table IV shows the natural time of program execution, and the Python script is executed on a MacBook Pro with 1.4GHz quad-core intel core i5 and 16GB 2133 MHz LPDDR3. For the HML meta-training, we execute $2,000$ training epochs and provide the average computation time of one epoch. For the HML inference and iterative algorithm, we calculate “sample 1” shown in Fig. 4(a) for $2000$ times, and provide the average computation time of this sample. We can observe that the time required for HML inference is significantly less than the HML meta-training. This is because meta-training involves extensive optimization across large datasets, adjusting parameters to learn patterns by backpropagation, but inference simply applies the learned parameters to make predictions on new data, which requires much lighter computations. We can also observe from this table that our HML reduces the computation time by approximately $200$ to $2000$ than the optimal iterative algorithm. Interestingly, in the iterative algorithm, the computation times for secrecy rates are consistently shorter than those for data rates. This is primarily due to the fewer number of users scheduled in the secrecy rates-related scenarios depicted in Fig. 5 compared to the data rates-related scenarios shown in Fig. 6, leading fewer iterations required for calculating the secrecy rates.
@@ VI Conclusion
In this paper, we developed an HML approach to train a GNN-based scalable bandwidth allocation policy that can generalize well in various communication scenarios, including different number of users, wireless channels, QoS requirements, and bandwidth. The main idea is to train the initial parameters of the GNN with various tasks in meta-training, and then fine-tune the parameters with a few samples in meta-testing. Simulation results showed that the performance gap between the GNN and the optimal policy obtained by an iterative algorithm is less than 55% in most of the cases. For unseen communication scenarios, the GNN can converge in $10$ to $30$ training epochs, which are much faster than the existing benchmarks. Numerical results validate that our HML can reduce the computation time by $200$ to $2000$ times compared with the optimal iterative algorithm. Our approach can be extended beyond bandwidth allocation, such as power allocation, precoding, and repetitions. Nevertheless, the feature engineering and the structure of GNN in other scenarios deserve further investigation.
@@ Appendix A Proof of Concavity for Secrecy Rate in Long Blocklength Regimes
$\begin{split}\frac{\partial r_{k}^{S,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})}{\partial{w_{\tau,k}^{D,\mathcal{I}}}}=&\frac{\partial\left(r_{k}^{D,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})-r_{k}^{e,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})\right)}{\partial w_{k}}\\ =&\frac{-\zeta_{k}+\ln\left(1+\frac{\zeta_{k}}{w_{\tau,k}^{D,\mathcal{I}}}\right)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})}\\ &-\frac{-\zeta_{k}^{e}+\ln\left(1+\frac{\zeta_{k}^{e}}{w_{\tau,k}^{D,\mathcal{I}}}\right)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})}\\ =&\frac{\ln\left(\dfrac{w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}}{w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e}}\right)}{\ln(2)}\\ &+\frac{(\zeta_{k}^{e}-\zeta_{k})w_{\tau,k}^{D,\mathcal{I}}}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})},\end{split}$ (A.1)
@@
$\begin{split}\frac{\partial^{2}r_{k}^{S,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})}{\partial{w_{\tau,k}^{D,\mathcal{I}}}^{2}}=&\frac{\partial}{\partial w_{\tau,k}^{D,\mathcal{I}}}\left(\frac{\partial r_{k}^{S,\mathcal{I}}(w_{\tau,k}^{D,\mathcal{I}})}{\partial{w_{\tau,k}^{D,\mathcal{I}}}}\right)\\ =&\dfrac{(\zeta_{k}^{e}-\zeta_{k})\left((\zeta_{k}^{e}+\zeta_{k})w_{\tau,k}^{D,\mathcal{I}}+2\zeta_{k}^{e}\zeta_{k}\right)}{\ln(2)(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k})^{2}(w_{\tau,k}^{D,\mathcal{I}}+\zeta_{k}^{e})^{2}}.\end{split}$ (A.2)